The athletic department of a school has 12 full time coaches and 4 of them are female. The director selects two coaches at random from this group. The probability that neither of them is a female is:
Answer by stanbon(75887)

The athletic department of a school has 12 full time coaches and 4 of them are female. The director selects two coaches at random from this group.
# of female 4 ; # of male 8
------
The probability that neither of them is a female is:
# of non-female pairs:: 8C2 = (8*7)/(1*2) = 28
# of random pairs:: 12C2 = (12*11)/(1*2) = 66
-----
Ans: P(2 non-female selected) = 28/66 = 42%
------
Cheers,
Stan H.
-----------